OpenStudy (matt6288):

CaH2(s) + 2 H2O(l)--->Ca(OH) 2(s) + 2 H2(g) After the reaction above, 90.0 mL of H2(g) was collected over water at 25 °C and a total pressure of 755.0 mm Hg. How many grams of CaH2 decomposed? The vapor pressure of H2O at 25 °C is 23.8 mm Hg?

OpenStudy (matt6288):

@sweetburger

OpenStudy (matt6288):

P:.9621atm V:.09L n:m/M R:.0821 L x atm/K x mol T:298.15K ok so I have to find grams of NaH2 so I made n=mass/molar mass m=MPV/RT and got .08847....g NaH2 this feel and looks wrong..

OpenStudy (matt6288):

@Photon336

OpenStudy (matt6288):

NaH2:25.01g

OpenStudy (matt6288):

using what numbers?

OpenStudy (matt6288):

RT/PV= how did you get .9934atm and 9x10^-2?

OpenStudy (photon336):

we have 90 ml right?

OpenStudy (matt6288):

90.0ml x 1L/1000mL=.09L

OpenStudy (photon336):

yep. and I just wrote that in scientific notation. \[9.0*10^{-2} L \]

OpenStudy (matt6288):

but i guess it doesnt matter how you write it

OpenStudy (matt6288):

how about .9934?

OpenStudy (matt6288):

what about the partial pressure? 755-23.8?

OpenStudy (photon336):

that's interesting. I'm thinking that we would need the total pressure for this

OpenStudy (photon336):

what did you have in mind?

OpenStudy (matt6288):

n=PV/RT .9621x.09/298.15x.0821=.003537 mols of H2

OpenStudy (matt6288):

right? .003537 H2 x 1molCaH2/2molsH2 x 25.01gCaH2/1molCaH2

OpenStudy (matt6288):

I used stoichiometry to find grams of CaH2

OpenStudy (matt6288):

I got .044...

OpenStudy (matt6288):

I give up >.<

OpenStudy (photon336):

let's start with the formula again PV = nRT \[ \frac{ PV }{ RT } = n \]

OpenStudy (photon336):

\[\frac{ \frac{ 731.2~mmHG }{ 760 } *(\frac{ 90 }{ 1000 }L )}{ 0.08*298K } = moles~H_{2}\]

OpenStudy (photon336):

\[3.63*10^{-3}~mol~H_{2}*(\frac{ CaH_{2} }{ 2H_{2}})*(\frac{ 42.094~Grams }{ Mole~CaH_{2} }) = 7.64*10^{-2}~grams \] 0.076 grams

OpenStudy (matt6288):

wait wheres ur R?

OpenStudy (matt6288):

you put .08 not .0821

OpenStudy (matt6288):

omg I made a mistake

OpenStudy (photon336):

I just rounded it up to 0.08

OpenStudy (matt6288):

I put 25.01g rather than 42.1g

OpenStudy (photon336):

yeah your answer is roughly half of mine

OpenStudy (matt6288):

so 0.076 is correct?

OpenStudy (photon336):

yeah, this is under the assumption that we subtracted the pressure of water from the total pressure.

OpenStudy (matt6288):

ya I kept thinking it was NaH2 but its CaH2 ;P

OpenStudy (matt6288):

thanks man your the best

OpenStudy (photon336):

yeah no problem. you knew what you were doing